OpenStudy (anonymous):

y varies directly as the square of x. when x=5, y=5. find y when x=10

OpenStudy (anonymous):

y=10 y=20 y=50 y=100

OpenStudy (aroub):

Directly! I thought inversely lol sorry... y=k(x^2) 5=k(5^2) 5=k(25) k=1/5 y=k(x^2) y=1/5(10^2) y=1/5(100) y=20 =D